OpenStudy (anonymous):

i will give a medal to best answer/become a fan. Find the solutions to the system. y=x^2+4x+6 y=4x+12

OpenStudy (anonymous):

replace \(y\) in the first equation by \(4x+12\) and solve \[4x+12=x^2+4x+6\]

OpenStudy (anonymous):

equate bothe the equation and solve for x

OpenStudy (anonymous):

\[4x+12=x ^{2}+4x+6 = x ^{2}-6\] \[x=\pm \sqrt{6}\]